OpenStudy (anonymous):

solve system of equation r-s=-3 r+s+9

OpenStudy (anonymous):

Lets start by separating the terms and work them out in groups. The first one we will do is: 3(r + 4) When you are multiplying by numbers or variable in parentheses, you distribute the number into the parentheses like this: (3*r + 3*4) this simplifies to: 3r + 12 next 2(r - 3) distribute like i did earlier: (2*r - 2*3) simplify: 2r - 6 and finally, the last one on the other side of the equation: 2(2r - 3) distribute again: 2*2r - 2*3 simplify: 4r - 6 So lets put Humpty back together again and we get: 3r + 12 + 2r - 6 = 4r - 6 remember that subtracting a number is the same as adding the complete opposite of that number, so now our equation becomes: 3r + 12 + 2r + (-6) = 4r + (-6) Now we combine like terms: 3r + 12 + 2r + (-6) = 4r + (-6) 5r + 12 + (-6) = 4r + (-6) 5r + 6 = 4r + (-6) now lets subtract 4r from both sides of the equation: 5r + 6 - 4r = 4r + (-6) - 4r r + 6 = (-6) Note that the 4r's on the right side of the equation cancel each other out. Now lets subtract 6 from both sides of the equation: r + 6 - 6 = (-6) - 6 r = (-6) - 6 r = (-12) Hope that helps :-)

jimthompson5910 (jim_thompson5910):

Add the two equations to get r-s=-3 r+s=9 ------- 2r+0s = 6 So 2r = 6 Solve that for r Then use the solution for r to find the solution for s